# Fading in the Flow: Suppression of cold gas growth in expanding galactic outflows

### v2

@@ Abstract
Multiwavelength observations reveal multiphase outflows that play a crucial role in redistributing gas and metals in and around galaxies. Theoretical modelling of such multiphase outflows often employs wind tunnel simulations of a spherical cold ($\sim 10^{4}\ \rm K$) cloud facing a uniform hot ($\sim 10^{6}\ \rm K$) wind. However, outflows are naturally expanding and wind conditions change downstream – a crucial aspect overlooked in most idealized simulations. To address this, we examine how an expanding wind influences the survival, morphology, and dynamics of a cloud. We perform idealized hydrodynamic simulations of radiative cloud-crushing in an expanding wind, where the steady background wind is modelled using the adiabatic Chevalier & Clegg 1985 (CC85) analytic solution. Moving downstream, we find that the clouds remain locally isobaric with the wind, leading to a steep decline in their density contrast with respect to the ambient medium, and they eventually dissipate into the wind. This also suppresses the growth of cold gas mass in comparison to a plane-parallel wind since entrained clouds move into a less radiative background. Using analytic scaling arguments, we present a physical picture of cloud evolution in a CC85 wind. Cloud expansion and local pressure equilibrium are the key regulators of cold mass growth. Unlike traditional homogeneous wind tunnel simulations, our simulations account for the differential expansion experienced by the long cometary tails of clouds moving in an outflow. Consequently, a strong head-to-tail emission gradient in the filamentary cold gas tails develop – features closer to observations. In addition, we demonstrate that the dynamics of individual clouds may substantially alter the radial properties of their host multiphase outflows.
@@ 1 Introduction
Galactic outflows, particularly those driven by starburst activities (Heckman et al. 1990; Lehnert & Heckman 1996), are pivotal in regulating galaxy evolution by expelling gas, metals, and energy (see Veilleux et al. 2020; Thompson & Heckman 2024 for reviews on galactic winds and Sarkar 2024 for a review on Fermi/eROSITA bubbles driven perpendicular to the Milky Way disk) into the circumgalactic medium (CGM; see Tumlinson et al. 2017; Faucher-Giguère & Oh 2023 for reviews on CGM) and even the intergalactic space in the case of stronger outflows, possibly powered by AGN (Oppenheimer et al. 2020; Davies et al. 2020; Keller et al. 2020). These outflows are inherently multiphase, encompassing hot, ionized gas, warm neutral and ionized components, and cold molecular clouds (Heckman et al. 2000; Thompson et al. 2016; Roberts-Borsani 2020; Veilleux et al. 2020; Vijayan et al. 2020; Fluetsch et al. 2021). Understanding the dynamics of these cold clouds within outflows is crucial, as they not only transport mass and metals, shaping star formation rates and influencing the thermal and chemical state of the CGM and the intergalactic medium (IGM), but also alter the overall dynamics of the wind itself through momentum and energy exchange (Hopkins et al. 2012; Anglés-Alcázar et al. 2017; Kim & Ostriker 2018; McQuinn & Werk 2018; Zabl et al. 2020; Belli et al. 2024).
@@
The interaction between cold clouds and hot winds is a fundamental process not only in galactic outflows but also in many other astrophysical contexts that influence galaxy formation and evolution (Dekel & Silk 1986; Cole et al. 1994; Erb 2008; Putman et al. 2012; Pezzulli & Fraternali 2016; Naab & Ostriker 2017). These include fountain flows (Shapiro & Field 1976; Bregman 1980; Houck & Bregman 1990; Fraternali & Binney 2008; Rubin et al. 2022; Marasco et al. 2022; Li et al. 2023) that potentially give rise to high-velocity clouds (Benjamin & Danly 1997; Putman 1999; Wakker 2001; Putman et al. 2002; Binney et al. 2009; Heitsch & Putman 2009; Richter et al. 2017; Lehner et al. 2022), and the stripping of gas from satellite/jellyfish galaxies (Roediger & Brüggen 2007; Tonnesen & Bryan 2010; Steinhauser et al. 2016; Ramatsoku et al. 2020; Franchetto et al. 2020; Poggianti et al. 2019; Campitiello et al. 2021; Deb et al. 2022) with the subsequent formation of cold gas (Ghosh et al. 2024; Roy et al. 2024) and stars (Tonnesen & Bryan 2012; Vulcani et al. 2018; Roberts et al. 2022) in their tails. The fate of these cold clouds, i.e., whether they survive and grow or are destroyed, has significant implications for the star formation rate/history of their host galaxies, as well as for the mass and energy budget of galaxies and their surrounding media. The cold gas may also condense from the hot and diffuse CGM due to thermal instability, particularly in the presence of large density fluctuations and stratification – common in CGM-like environments (e.g., Sharma et al. 2012; McCourt et al. 2012; Voit 2018; Choudhury et al. 2019; Mohapatra et al. 2020, 2022; Wibking et al. 2025). Most importantly, observations indicate the presence of multiphase gas for most sightlines through the CGM, indicating the ubiquity of cold and hot gas sitting next to each other in different parts of galactic atmospheres.
@@
The importance of cold clouds interacting with the surrounding medium in shaping galaxy formation has also been recently demonstrated in cosmological simulations of galaxy formation (Nelson et al. 2020; Ramesh & Nelson 2024). Observations as well as simulations suggest that similar physics of multiphase gas mixing governs the formation of Magellanic streams surrounding the large and small Magellanic clouds (D’Onghia & Fox 2016; Fox et al. 2020; Lucchini et al. 2021; Mishra et al. 2024) as well as in the fueling galaxies with cold gas by pristine cold filaments from the IGM (Kereš et al. 2005; Dekel et al. 2009; van de Voort et al. 2012; Mandelker et al. 2016, 2020; Berlok & Pfrommer 2019; Martin et al. 2019; Aung et al. 2024), especially around the cosmic noon ($z\sim 2-3$). Even seemingly disparate astrophysical phenomena in terms of time and length scales involved, such as the Coronal rain in the Sun, exhibit astonishing similarities with cloud-wind interaction, multiphase gas mixing, and thermal instability occurring in galactic atmospheres (Antolin et al. 2010; Antolin 2020; Kolotkov et al. 2023; Hillier et al. 2025).
@@
These early ‘cloud crushing’ studies have revealed a common picture in which cold clouds undergo mixing and destruction due to a combination of various hydrodynamic instabilities, such as Kelvin-Helmholtz, Rayleigh-Taylor, and Richtmyer-Meshkov, which develop as a consequence of the relative motion and subsequent momentum transfer between the cloud and the wind. A critical timescale in this context is the cloud-crushing time $t_{\rm cc}$, defined as the timescale over which a shock propagates through the cloud, leading to its compression and potential destruction. This timescale, which also approximately equals the time for the growth of the unstable modes of the Kelvin-Helmholtz instability, is given by $t_{\rm cc}=\sqrt{\chi}R_{\rm cl}/v_{\rm wind}$, where $\chi$ is the factor by which the cloud is over-dense compared to the wind, $R_{\rm cl}$ is the cloud size, and $v_{\rm wind}$ is the wind velocity.
@@
A significant breakthrough in the relevance of the cloud-crushing problem to the multiphase CGM came with the discovery of a regime in which cold clouds can grow in the wind in the presence of radiative cooling (Armillotta et al. 2016, 2017; Gronke & Oh 2018; Li et al. 2020; Sparre et al. 2020; Kanjilal et al. 2021). In order to explain the origin of high-velocity clouds (HVCs) and cold extra-planar HI gas in Milky Way-like galaxies (Oosterloo et al. 2007; Fraternali 2009), Marinacci et al. 2010, 2011 carried out two-dimensional simulations of clouds (possibly uplifted from the ISM by supernovae) falling towards the Galactic disk. These are some of the earliest cloud-crushing simulations that found growth of cold gas due to radiative cooling in the wakes of clouds moving in the CGM (see review by Fraternali 2017). More recent simulations have supported this finding and have consistently shown that clouds can grow in the regime of rapid radiative cooling. The relative shear between the cold cloud and the hot wind gives rise to a turbulent boundary layer with a short cooling time. This can lead to the entrainment of the diffuse gas into the dense medium. Such turbulent radiative mixing layers (TRMLs) have been extensively studied in recent idealized simulations covering a wide range of astrophysical conditions (Ji et al. 2019; Fielding et al. 2020; Tan et al. 2021; Yang & Ji 2023; Abruzzo et al. 2024), providing a general consistency with the results of radiative cloud-crushing simulations. While TRMLs driven by shear lead to mass growth at early times in radiative cloud crushing, after entrainment, the relative velocity between the cloud and the wind decreases and yet the dense cloud growth rate peaks in such a state without shear (Gronke & Oh 2020). This late-time growth may be interpreted as the steady, subsonic cooling flow solution that exists in spherical and cylindrical flow geometries (no such steady inflow solution exists in Cartesian geometry; Dutta et al. 2022).
@@
$\approx 10\ {\rm pc}\left(\frac{T_{\rm cl}}{10^{4}\rm\ K}\right)^{5/2}\left(\frac{p/k_{B}}{10^{3}\ \rm K\ cm^{-3}}\right)^{-1}\left(\frac{\mathcal{M}}{1.0}\right)$
@@
However, real galactic outflows are more complex, often exhibiting adiabatic expansion as they move downstream due to the continuous energy and mass input from starburst regions (Chevalier & Clegg 1985). This expansion leads to spatial and temporal variations in wind properties that the cloud experiences as it moves down the wind, necessitating a more generalized approach to studying cloud-wind interactions. In this paper, we extend the traditional cloud-crushing model. Moving beyond the traditional model where the background wind was considered to be uniform (which we refer to as ‘vanilla’ cloud-crushing), we investigate the behavior of cold clouds in adiabatically expanding winds powered by continuous starbursts. Through a series of hydrodynamic simulations, we analyze how these realistic wind environments influence the survival, acceleration, and mixing of clouds. We model the radially expanding adiabatic background wind using the steady-state analytic solution of an adiabatically expanding spherical wind proposed by Chevalier & Clegg 1985 (CC85 hereafter) for starburst-driven outflows. In a CC85 wind, an injection radius exists within which continuous star formation (and subsequent supernovae) injects mass and (thermal-)energy at a constant rate, which drives a steady wind. The wind transitions from a subsonic regime to a supersonic regime as it moves past the injection radius, where the Mach number of the wind is unity. In the asymptotic limit, the CC85 profile gives the wind pressure $p_{\rm wind}\propto d^{-2\gamma}$, density $\rho_{\rm wind}\propto d^{-2}$, and radial velocity $v_{\rm wind}\sim\rm{constant}$, where $d$ is the distance from the center of the wind and $\gamma$ is the ratio of specific heats.
Our paper is organized as follows. We present our numerical setup and discuss some of the novel techniques developed for our study in Section 2. In Section 3, we present our simulation results and analyze the cloud properties as they evolve in a CC85 wind. We show that the growth of cold clouds is suppressed in an expanding outflow and clouds become increasingly indistinguishable from the background wind as they travel downstream. In Section 4, we compare with previous simulations, and models of cloud-wind interaction and multiphase galactic outflows. We discuss models that evolve the clouds in an expanding wind using the analytic framework of radiatively cooling turbulent boundary layers. We build a physical theory of cloud evolution based on its downstream expansion, assuming it to be locally isobaric relation with the wind, backed by our simulation results. In this context, we also highlight the limitations of the existing models and simulations of cloud-wind interaction and how model parameters dramatically affect the properties of multiphase galactic outflows. We also compare with recent JWST observations of cold clouds in outflows (in M82). Finally, we summarize and conclude in Section 5.
@@ 2 Physical setup & Numerical Simulations
Our numerical simulations use a three-dimensional grid with spherical polar coordinates. Our grid is uniformly spaced along the radial as well as along the polar and azimuthal directions. The spherical polar geometry of the grid aligns with the spherical symmetry of the CC85 wind and reduces numerical artifacts that can otherwise perturb the steady-state CC85 wind profile used as the background. We have used a modified version of PLUTO v4.4p2 in our simulations. It is a conservative finite-volume Godunov (magneto) hydrodynamics code (Mignone et al. 2007). We also include optically thin radiative cooling of gas at solar metallicity assuming photo+collisional ionization equilibrium in the presence of a ($z=0$) Haardt-Madau extragalactic UV background (Haardt & Madau 2012). Radiative cooling in our simulation is turned off at the floor temperature of $4\times 10^{4}$ K, the initial temperature of the cloud. This mimics the drastic reduction in cooling efficiency below this temperature because of the heating by the extragalactic background radiation. Further, in order to prevent deviations from the steady adiabatically expanding background, we have turned off radiative cooling in the bulk of the wind (identified with passive tracer), and only the mixed gas is allowed to cool radiatively. Thermal instability in the bulk of the background wind in the presence of radiative cooling (Wang 1995; Thompson et al. 2016; Schneider et al. 2018; Nguyen et al. 2024) is a related but different problem – beyond the scope of our present study. We have also not considered other physical processes like magnetic fields, thermal conduction, self-gravity, etc. Since this is the first work that investigates the effect of an expanding background on cloud-crushing simulations, we restrict ourselves to the minimum set of physical effects. In this controlled simulation setup, we explore the parameter space of cloud-wind interaction in an expanding background and prepare the necessary groundwork before introducing additional physical effects in the future. An additional caveat with the choice of spherical coordinates is that the cell size increases worsensing the resolution downstream. However, on comparing cold mass across simulations at different resolutions (cf. Appendix D), we find the mass estimate agrees within a factor of a few – enough for our simulations to provde a robust estimate of the different scaling arguments (presented later in the paper) and reliably capture the physics of cloud-wind interaction.
2.1 Initial & boundary conditions
Our initial condition is a stationary, dense, spherical cloud of radius $R_{\rm cl}$ at a temperature of $T_{\rm cl}=4\times 10^{4}\ \rm K$ in the simulation domain. The cloud is initialized to be locally isobaric with the background wind. The density of the cloud (hence its temperature) is initialized to a fixed contrast $\chi$ with the background wind. The background condition outside the cloud is determined by the steady spherically symmetric solution of a CC85 wind (Chevalier & Clegg 1985). Furthermore, we have ensured that our computational domain always excludes the region within the injection radius (sonic point) of the CC85 wind. This simplifies our numerical scheme, as the fluid equations beyond the sonic point do not require the inclusion of any mass, momentum, or energy injection sources. We were able to maintain a stable and steady background wind profile in the entire simulation domain consistent with the CC85 solution by fixing the boundary condition to any particular CC85 wind profile.

### v3

@@ Abstract
Multiphase outflows, revealed by multi-wavelength observations, are crucial in redistributing gas and metals within and around galaxies. These outflows are often modelled theoretically using wind tunnel simulations of a cold ($\sim 10^{4}$ K) cloud interacting with a uniform hot ($\sim 10^{6}$ K) wind. However, real outflows expand downstream, a feature overlooked in most idealised simulations. We study how an expanding wind affects the survival, morphology, and dynamics of a cloud. We conduct idealised hydrodynamic simulations with optically thin radiative cooling of a cloud in an expanding wind, modelled using the adiabatic Chevalier & Clegg (1985) solution. We find that clouds remain locally isobaric with the wind, leading to a steep decline in their density contrast and eventual dissolution downstream. Compared to a plane-parallel wind, this suppresses cold gas mass growth because as clouds travel downstream, the surrounding mixed boundary layer becomes diffuse and less radiative. Our analytical scaling arguments show that cloud expansion and local pressure equilibrium are the key regulators of cold mass growth. Unlike traditional simulations, our model accounts for the differential expansion experienced by the long cometary tails of clouds in wind tunnels. This creates a strong head-to-tail emission gradient in the filamentary cold gas, which is more consistent with observations. We also demonstrate that the dynamics of individual clouds can substantially alter the radial properties of their host multiphase outflows.
@@ 1 Introduction
Galactic outflows, particularly those driven by starburst activities (Heckman et al. 1990; Lehnert & Heckman 1996), are pivotal in regulating galaxy evolution by expelling gas, metals, and energy (see Veilleux et al. 2020; Thompson & Heckman 2024 for reviews on galactic winds and Sarkar 2024 for a review on Fermi/eROSITA bubbles driven perpendicular to the Milky Way disk) into the circumgalactic medium (CGM; see Tumlinson et al. 2017; Faucher-Giguère & Oh 2023 for reviews on CGM) and even the intergalactic space in the case of stronger outflows, possibly powered by AGN (Oppenheimer et al. 2020; Davies et al. 2020; Keller et al. 2020). These outflows are inherently multiphase, encompassing hot, ionised gas, warm neutral and ionised components, and cold molecular clouds (Heckman et al. 2000; Thompson et al. 2016; Roberts-Borsani 2020; Veilleux et al. 2020; Vijayan et al. 2020; Fluetsch et al. 2021). Understanding the dynamics of these cold clouds within outflows is crucial, as they not only transport mass and metals, shaping star formation rates and influencing the thermal and chemical state of the CGM and the intergalactic medium (IGM), but also alter the overall dynamics of the wind itself through momentum and energy exchange (Hopkins et al. 2012; Anglés-Alcázar et al. 2017; Kim & Ostriker 2018; McQuinn & Werk 2018; Zabl et al. 2020; Belli et al. 2024).
@@
The interaction between cold clouds and hot winds is a fundamental process not only in galactic outflows but also in many other astrophysical contexts that influence galaxy formation and evolution (Dekel & Silk 1986; Cole et al. 1994; Erb 2008; Putman et al. 2012; Pezzulli & Fraternali 2016; Naab & Ostriker 2017). These include fountain flows (Shapiro & Field 1976; Bregman 1980; Houck & Bregman 1990; Fraternali & Binney 2008; Rubin et al. 2022; Marasco et al. 2022; Li et al. 2023) that potentially give rise to high-velocity clouds (Benjamin & Danly 1997; Putman 1999; Wakker 2001; Putman et al. 2002; Binney et al. 2009; Heitsch & Putman 2009; Richter et al. 2017; Lehner et al. 2022), and the stripping of gas from satellite/jellyfish galaxies (Roediger & Brüggen 2007; Tonnesen & Bryan 2010; Steinhauser et al. 2016; Ramatsoku et al. 2020; Franchetto et al. 2020; Poggianti et al. 2019; Campitiello et al. 2021; Deb et al. 2022) with the subsequent formation of cold gas (Ghosh et al. 2024; Roy et al. 2024) and stars (Tonnesen & Bryan 2012; Vulcani et al. 2018; Roberts et al. 2022) in their tails. The fate of these cold clouds, i.e., whether they survive and grow or are destroyed, has significant implications for the star formation rate/history of their host galaxies, as well as for the mass and energy budget of galaxies and their surrounding media. The cold gas may also condense from the hot and diffuse CGM due to thermal instability, particularly in the presence of large density fluctuations and stratification – common in CGM-like environments (e.g., Sharma et al. 2012; McCourt et al. 2012; Voit 2018; Choudhury et al. 2019; Mohapatra et al. 2020, 2022; Wibking et al. 2025). Most importantly, observations indicate the presence of multiphase gas for most sightlines through the CGM, indicating the ubiquity of cold and hot gas sitting next to each other in different parts of galactic atmospheres. At smaller scales - particularly when probing the structure and thermodynamics of multiphase gas in the interstellar medium (ISM) - numerous studies have explored the physics of cloud-wind interactions. These include interactions between clouds and supernova-driven winds within the ISM (Cowie et al., 1981; Pittard, 2019; Kim et al., 2017; Guo et al., 2024), as well as with the immediate surroundings of supernova remnants (Dwarkadas, 2007), and in bubbles inflated by stellar winds (Lancaster et al., 2021; Pittard, 2022; Dwarkadas, 2023). Remarkably, even seemingly unrelated astrophysical phenomena involving vastly different time and length scales - such as coronal rain in the solar atmosphere - exhibit striking similarities to cloud-wind interactions including multiphase gas mixing and thermal instability, also observed in galactic environments (Antolin et al., 2010; Antolin, 2020; Kolotkov et al., 2023; Hillier et al., 2025).
@@
The importance of cold clouds interacting with the surrounding medium in shaping galaxy formation has also been recently demonstrated in cosmological simulations of galaxy formation (Nelson et al. 2020; Ramesh & Nelson 2024). Observations as well as simulations suggest that similar physics of multiphase gas mixing governs the formation of Magellanic streams surrounding the large and small Magellanic clouds (D’Onghia & Fox 2016; Fox et al. 2020; Lucchini et al. 2021; Mishra et al. 2024) as well as in the fuelling galaxies with cold gas by pristine cold filaments from the IGM (Kereš et al. 2005; Dekel et al. 2009; van de Voort et al. 2012; Mandelker et al. 2016, 2020; Berlok & Pfrommer 2019; Martin et al. 2019; Aung et al. 2024), especially around the cosmic noon ($z\sim 2-3$).
@@
These early ‘cloud-crushing’ studies have revealed a common picture in which cold clouds undergo mixing and destruction due to a combination of various hydrodynamic instabilities, such as Kelvin-Helmholtz (von Helmholtz 1868; Thomson 1871; Chandrasekhar 1961), Rayleigh-Taylor (Rayleigh 1882; Taylor 1950; Kull 1991; Choudhuri 1998), and Richtmyer-Meshkov (Richtmyer 1960; Meshkov 1969; Yang et al. 1994; Brouillette 2002; Zhou et al. 2021; Napieralski et al. 2024), which develop as a consequence of the relative motion and subsequent momentum transfer between the cloud and the wind. A critical timescale in this context is the cloud-crushing time $t_{\rm cc}$, defined as the timescale over which a shock propagates through the cloud (see Eqs. 2.2 & 2.4 from Klein et al. 1994), leading to its compression and potential destruction of the cloud. This timescale, which also approximately equals the time for the growth of the unstable modes of the Kelvin-Helmholtz instability, is given by $t_{\rm cc}=\sqrt{\chi}R_{\rm cl}/v_{\rm wind}$, where $\chi$ is the factor by which the cloud is over-dense compared to the wind, $R_{\rm cl}$ is the cloud size, and $v_{\rm wind}$ is the wind velocity.
@@
A significant breakthrough in the relevance of the cloud-crushing problem to the multiphase CGM came with the discovery of a regime in which cold clouds can grow in the wind in the presence of radiative cooling (Armillotta et al. 2016, 2017; Gronke & Oh 2018; Li et al. 2020; Sparre et al. 2020; Kanjilal et al. 2021). In order to explain the origin of high-velocity clouds (HVCs) and cold extra-planar HI gas in Milky Way-like galaxies (Oosterloo et al. 2007; Fraternali 2008), Marinacci et al. 2010, 2011 carried out two-dimensional simulations of clouds (possibly uplifted from the ISM by supernovae) falling towards the Galactic disk. These are some of the earliest cloud-crushing simulations that found growth of cold gas due to radiative cooling in the wakes of clouds moving in the CGM (see review by Fraternali 2017). More recent simulations have supported this finding and have consistently shown that clouds can grow in the regime of rapid radiative cooling. The relative shear between the cold cloud and the hot wind gives rise to a turbulent boundary layer with a short cooling time. This can lead to the entrainment of the diffuse gas into the dense medium. Such turbulent radiative mixing layers (TRMLs) have been extensively studied in recent idealised simulations covering a wide range of astrophysical conditions (Ji et al. 2019; Fielding et al. 2020; Tan et al. 2021; Yang & Ji 2023; Abruzzo et al. 2024), providing a general consistency with the results of radiative cloud-crushing simulations. While TRMLs driven by shear lead to mass growth at early times in radiative cloud-crushing, after entrainment, the relative velocity between the cloud and the wind decreases and yet the dense cloud growth rate peaks in such a state without shear (Gronke & Oh 2020). This late-time growth may be interpreted as the steady, subsonic cooling flow solution that exists in spherical and cylindrical flow geometries (no such steady inflow solution exists in Cartesian geometry; Dutta et al. 2022).
@@
[…]
@@
However, real galactic outflows are more complex, often exhibiting adiabatic expansion as they move downstream due to the continuous energy and mass input from starburst regions (Chevalier & Clegg 1985). This expansion leads to spatial and temporal variations in wind properties that the cloud experiences as it moves down the wind, necessitating a more generalised approach to studying cloud-wind interactions. In this paper, we extend the traditional cloud-crushing model. Moving beyond the traditional model where the background wind was considered to be uniform (which we refer to as ‘vanilla’ cloud-crushing), we investigate the behaviour of cold clouds in adiabatically expanding winds powered by continuous starbursts. Through a series of hydrodynamic simulations, we analyse how these realistic wind environments influence the survival, acceleration, and mixing of clouds. We model the radially expanding adiabatic background wind using the steady-state analytic solution of an adiabatically expanding spherical wind proposed by Chevalier & Clegg 1985 (CC85 hereafter) for starburst-driven outflows. In a CC85 wind, an injection radius exists within which continuous star formation (and subsequent supernovae) injects mass and (thermal-)energy at a constant rate, which drives a steady wind. The wind transitions from a subsonic regime to a supersonic regime as it moves past the injection radius, where the Mach number of the wind is unity. In the asymptotic limit, the CC85 profile gives the wind pressure $p_{\rm wind}\propto d^{-2\gamma}$, density $\rho_{\rm wind}\propto d^{-2}$, and radial velocity $v_{\rm wind}\sim\rm{constant}$, where $d$ is the distance from the centre of the wind and $\gamma$ is the ratio of specific heats.
Our paper is organized as follows. We present our numerical setup and discuss some of the novel techniques developed for our study in Section 2. In Section 3, we present our simulation results and analyse the cloud properties as they evolve in a CC85 wind. We show that the growth of cold clouds is suppressed in an expanding outflow and clouds become increasingly indistinguishable from the background wind as they travel downstream. In Section 4, we compare with previous simulations, and models of cloud-wind interaction and multiphase galactic outflows. We discuss models that evolve the clouds in an expanding wind using the analytic framework of radiatively cooling turbulent boundary layers. We build a physical theory of cloud evolution based on its downstream expansion, assuming it to be locally isobaric relation with the wind, backed by our simulation results. In this context, we also highlight the limitations of the existing models and simulations of cloud-wind interaction and how model parameters dramatically affect the properties of multiphase galactic outflows. We also compare with recent JWST observations of cold clouds in outflows (in M82). Finally, we summarize and conclude in Section 5.
@@ 2 Physical setup & Numerical Simulations
Our numerical simulations use a three-dimensional grid with spherical polar coordinates. Our grid is uniformly spaced along the radial as well as along the polar and azimuthal directions. The spherical polar geometry of the grid aligns with the spherical symmetry of the CC85 wind and reduces numerical artifacts that can otherwise perturb the steady-state CC85 wind profile used as the background. We have used a modified version of PLUTO v4.4p2 in our simulations. It is a conservative finite-volume Godunov (magneto) hydrodynamics code (Mignone et al. 2007). We also include optically thin radiative cooling of gas at solar metallicity assuming photo+collisional ionisation equilibrium in the presence of a ($z=0$) Haardt-Madau extragalactic UV background (Haardt & Madau 2012). Radiative cooling in our simulation is turned off at the floor temperature of $4\times 10^{4}$ K, the initial temperature of the cloud. This mimics the drastic reduction in cooling efficiency below this temperature because of the heating by the extragalactic background radiation. Further, in order to prevent deviations from the steady adiabatically expanding background, we have turned off radiative cooling in the bulk of the wind (identified with passive tracer), and only the mixed gas is allowed to cool radiatively. Thermal instability in the bulk of the background wind in the presence of radiative cooling (Wang 1995; Thompson et al. 2016; Schneider et al. 2018; Nguyen et al. 2024) is a related but different problem – beyond the scope of our present study. We have also not considered other physical processes like magnetic fields, thermal conduction, self-gravity, etc. Since this is the first work that investigates the effect of an expanding background on cloud-crushing simulations, we restrict ourselves to the minimum set of physical effects. In this controlled simulation setup, we explore the parameter space of cloud-wind interaction in an expanding background and prepare the necessary groundwork before introducing additional physical effects in the future. An additional caveat with the choice of spherical coordinates is that the cell size increases, worsening the resolution downstream. However, on comparing cold mass across simulations at different resolutions (cf. Appendix E), we find the mass estimate agrees within a factor of a few – enough for our simulations to provide a robust estimate of the different scaling arguments (presented later in the paper) and reliably capture the physics of cloud-wind interaction.